The perimeter of the square is 20 cm. Find it’s area.
Include the units.

Answer by Theo(13342)

the equation for the perimeter of the square is p = 4s
the equation for the area of the square id a = s^2
when p = 20, the perimeter becomes 20 = 4s
solve for s to get s = 20/4 = 5
the area becomes a = 5^2 = 25.
the perimeter and side are in cm.
the area is in cm^2 (square centimeters).